# TESS Investigation - Demographics of Young Exoplanets (TI-DYE) II: a second giant planet in the 17-Myr system HIP 67522

### v1

@@ Abstract
The youngest (50 Myr) planets are vital to understand planet formation and early evolution. The 17 Myr system HIP 67522 is already known to host a giant (10) planet on a tight orbit. In the discovery paper, Rizzuto et al. 2020 reported a tentative single transit detection of an additional planet in the system using TESS. Here, we report the discovery of HIP 67522 c which matches with that single transit event. We confirm the signal with ground-based multi-wavelength photometry from Sinistro and MuSCAT4. At a period of 14.33 days, planet c is close to a 2:1 mean motion resonance with b (6.96 days or 2.06:1). The light curve shows distortions during many of the transits, which are consistent with spot crossing events and/or flares. Fewer stellar activity events are seen in the transits of planet b, suggesting that planet c is crossing a more active latitude. Such distortions, combined with systematics in the TESS light curve extraction, likely explain why planet c was previously missed.
I Introduction
@@
Young ( Myr) planets provide an opportunity to detect planet-sculpting processes as they happen. In particular, young multi-planet systems are valuable because they allow for better control of variables within a single system and enable a wider range of science cases than those with single planets. Such systems enable tests on the origin of intrasystem uniformity (e.g., Lissauer et al., 2011; Lammers et al., 2023), measure more precise planetary eccentricities (e.g., Van Eylen et al., 2019), and search for transit timing variations that may yield precise masses and eccentricities (e.g. Steffen et al., 2012; Masuda, 2014), among a wide range of other science cases.
@@
Stellar, and hence planetary, ages are generally challenging to determine for individual systems (Soderblom et al., 2014). Therefore, planets in stellar associations are critical in understanding planet formation mechanisms, as the planets’ ages can be precisely derived from the parent stellar population. Studies from the small sample of young planets (30) have demonstrated several key findings: some close-in planets either form in situ or migrate quickly (David et al., 2019a; Mann et al., 2022; Wood et al., 2023), young planets tend to be larger than their older counterparts (Fernandes et al., 2022, 2023; Vach et al., 2024), and planets generally form and/or migrate while maintaining alignment with their host star (Hirano et al., 2020; Johnson et al., 2022; Hirano et al., 2024).
@@
TESS has played a critical role in the discovery and characterization of young planetary systems. TESS data was used to discover systems as young as 11 Myr (Mann et al., 2022; Zakhozhay et al., 2022) and young planets as small as (Livingston et al., 2018; Capistrant et al., 2024), as well as aided in the discovery of new young stellar associations hosting transiting planets (e.g, Tofflemire et al., 2021; Nardiello et al., 2022; Thao et al., 2024).
@@
II TESS light curve
@@
II.1 Extraction pipeline
For our analysis, we used a custom light curve extraction pipeline starting from Simple Aperture Photometry (SAP; Twicken et al., 2010; Morris et al., 2020) curves produced by the Science Processing Operations Center (SPOC; Jenkins et al., 2016). We used the shortest cadence data available for each sector.
We applied systematic corrections following the prescription in Vanderburg et al. (2019). To summarize, we corrected the SPOC SAP light curves with a linear model consisting of a basis spline (B-spline) with regularly-spaced breaks at 0.2 day intervals to model long-term, low-frequency stellar variability, several moments of the distribution of the spacecraft quaternion time series measurements within each exposure, seven co-trending basis vectors from the SPOC Presearch Data Conditioning (PDC; Stumpe et al., 2012; Smith et al., 2012; Stumpe et al., 2014) band-3 flux time series correction with the largest eigenvalues, and a high-pass (0.1 day) time series from the SPOC background aperture.
@@ II.1 Extraction pipeline
III Identification of the transit
As part of our search for transiting planets in the youngest stellar associations (Barber et al. submitted), and motivated by both the single-transit detection in Rizzuto et al. (2020) and the recent transit timing variations (TTVs) detected in HIP 67522 b (Thao et al. submitted), we searched HIP 67522 for additional planet signals. We used the updated N&L pipeline (Rizzuto et al., 2017) as described in Barber et al. (submitted). To summarize, we updated the default box-least squares (BLS) search to the more recent implementation in astropy111http://www.astropy.org (version 4.2; Astropy Collaboration et al., 2018) and used a more optimized period and duration search grid.
@@ III Identification of the transit
IV Ground-based follow up
IV.1 MuSCAT4
@@ IV.2 Sinistro
V Planet properties
V.1 TESS transit fit
We fit the systematics-corrected TESS photometry using MISTTBORN (MCMC Interface for Synthesis of Transits, Tomography, Binaries, and Others of a Relevant Nature; Mann et al., 2016; Johnson et al., 2018), which uses BATMAN to generate model transits (Kreidberg, 2015), celerite to model stellar variability with a Gaussian Process (GP; Foreman-Mackey et al., 2017), and emcee to explore the parameter space (Foreman-Mackey et al., 2013).
We fit both planets (b and c) and the stellar variability simultaneously, with 20 fit parameters in total. For each planet, we fit for the time of periastron (), planet orbital period (), planet-to-star radius ratio (), impact parameter (), and and to account for orbital eccentricity () and the argument of periastron (). We fit both planets with a common stellar density () and triangular limb-darkening coefficients (, ; Kipping, 2013).
The remaining five parameters were used to fit for the stellar variability in the GP model. We used two stochastically-driven damped simple harmonic oscillators (SHOMixture), one fixed to be at half the period of the other (Foreman-Mackey et al., 2017). The free parameters were the GP period (), the GP amplitude (), the quality factor for the simple harmonic oscillators (described by and ), and a mixture term (Mix).
Most parameters evolved under uniform priors with only physical limitations. For and , we found that 1-2 (of 50) walkers would wander off the true transit time, preferring to fit nearby stellar variations or flares as one of the transits. To fix this, we limited both and to within 0.01 days (15 minutes) of the discovery values for c and the values from Rizzuto et al. (2020) for b. This limit is small compared to the final uncertainties and lower than the TTV amplitude (Thao et al. submitted). Visual inspection and comparison to the final parameters suggest this did not introduce any bias (the final uncertainties were much smaller). For , we initialized the value and prior following the stellar parameters in Rizzuto et al. (2020) (/ = ). Using the LDTK toolkit (Parviainen & Aigrain, 2015), we initialized the limb-darkening coefficients ( and ).
@@ V.1 TESS transit fit
V.2 Ground-based transit fits
@@
In the MuSCAT4 data, there is a perturbation between 2460422.985 and 2460423.05 BJD. Similarly, in the Sinistro simultaneous transits, all three datasets exhibit a small spike in flux around 2460437.445 BJD. As we discuss further in Section VII, these were most likely spot crossings and were also seen in the TESS data. Some events may be flares (particularly the first bump in the MuSCAT4 data), however, our Gaussian spot model (below) described the deviations well.
@@
VI False positive analysis
@@
VII Summary and Conclusions
We report the discovery and validation of HIP 67522 c, originally identified as a single transit event by Rizzuto et al. (2020). We find HIP 67522 c to be 0.73 (8.2 ) and near the 2:1 mean motion resonance (MMR) with HIP 67522 b. We confirm the period and transit depth with two follow-up ground-based transit observations taken with MuSCAT4 and Sinistro. The HIP 67522 system is now the youngest known transiting multi-planet system to-date.
@@ VII Summary and Conclusions
Many of the young transiting multi-planet systems are also near MMR (e.g. AU Mic b & c (9:4) (Plavchan et al., 2020; Martioli et al., 2021), V1289 Tau c & d (3:2) and d & b (2:1) (David et al., 2019b, a; Feinstein et al., 2022), HD 109833 b & c (3:2) (Wood et al., 2023), and TIC 434398831 b & c (5:3) (Vach et al., 2024, in prep). Simulations suggest that planets will form in MMR and will begin to drift once the disk dissipates (5 Myr) (Izidoro et al., 2017), although many mature planets are found at or near MMR (e.g. Steffen & Hwang, 2015). Hamer & Schlaufman (2024) suggest that MMRs should be most common in systems from 10-100 Myr, and Dai et al. (2024) note that such an excess is visible in the known population of young planets.
@@
HIP 67522 b was a Cycle 1 JWST target (Mann et al., 2021). The resulting transmission spectrum showed strong features, with the transit depth varying by 50% in the CO2 and H2O bands, consistent with a low-density planet (Thao et al. submitted). HIP 67522 c is only 20% smaller than HIP 67522 b (0.73 vs 0.90). Assuming it also has a low density, as predicted for such young planets (Owen, 2020), it will be a similarly high-priority target for JWST.
@@ Acknowledgments
The authors wish to thank Halee, Wally, Penny, and Bandit for their invaluable support. M.G.B. was supported by NSF Graduate Research Fellowship (DGE-2040435), the NC Space Grant Graduate Research Fellowship, and the TESS Guest Investigator Cycle 5 program (21-TESS21-0016; NASA Grant number 80NSSC24K0880). P.C.T. was supported by NSF Graduate Research Fellowship (DGE-1650116), the Jack Kent Cooke Foundation Graduate Scholarship, and the JWST GO program (2498). A.W.M. was supported by the NSF CAREER program (AST-2143763) and a grant from NASA’s Exoplanet Reseach Program (XRP 80NSSC21K0393). This work is partly supported by JSPS KAKENHI Grant Number JPJP24H00017 and JSPS Bilateral Program Number JPJSBP120249910.

### v2

@@ Abstract
The youngest (50 Myr) planets are vital to understand planet formation and early evolution. The 17 Myr system HIP 67522 is already known to host a giant (10) planet on a tight orbit. In the discovery paper, Rizzuto et al. 2020 reported a tentative single transit detection of an additional planet in the system using TESS. Here, we report the discovery of HIP 67522 c, a 7.9 planet which matches with that single transit event. We confirm the signal with ground-based multi-wavelength photometry from Sinistro and MuSCAT4. At a period of 14.33 days, planet c is close to a 2:1 mean motion resonance with b (6.96 days or 2.06:1). The light curve shows distortions during many of the transits, which are consistent with spot crossing events and/or flares. Fewer stellar activity events are seen in the transits of planet b, suggesting that planet c is crossing a more active latitude. Such distortions, combined with systematics in the TESS light curve extraction, likely explain why planet c was previously missed.
1 Introduction
@@
Young ( Myr) planets provide an opportunity to detect planet-sculpting processes as they happen. In particular, young multi-planet systems are valuable because they allow for better control of variables within a single system and enable a wider range of science cases than those with single planets. Such systems enable tests on the origin of intrasystem uniformity (e.g., Lissauer et al., 2011; Lammers et al., 2023), differential measurements of planetary atmsopheres within a system (comparative planetology, e.g., Barat et al., 2024), and searches for transit timing variations that may yield precise masses and eccentricities (e.g. Steffen et al., 2012; Masuda, 2014), among a wide range of other science cases.
@@
Stellar, and hence planetary, ages are generally challenging to determine for individual systems (Soderblom et al., 2014). Therefore, planets in stellar associations are critical in understanding planet formation mechanisms, as the planets’ ages can be precisely derived from the parent stellar population. Studies from the small sample of young planets (30) have demonstrated several key findings: some close-in planets either form in situ or migrate quickly (David et al., 2019a; Mann et al., 2022; Wood et al., 2023), young planets tend to be larger than their older counterparts (Fernandes et al., 2022, 2023; Vach et al., 2024a), and among the six Myr planets with Rossiter–McLaughlin or Doppler tomography, all are aligned or nearly aligned (e.g. Hirano et al., 2020; Johnson et al., 2022; Hirano et al., 2024).
@@
TESS (Ricker et al., 2015) and K2 (Howell et al., 2014) have played critical roles in the discovery and characterization of young planetary systems. Light curves from these missions have been used to discover systems as young as 11 Myr (Mann et al., 2022; Zakhozhay et al., 2022) and young planets as small as (Livingston et al., 2018; Capistrant et al., 2024), as well as aided in the discovery of new young stellar associations hosting transiting planets (e.g, Tofflemire et al., 2021; Nardiello et al., 2022; Thao et al., 2024).
@@
2 TESS light curve
@@
2.1 Extraction pipeline
For our analysis, we used a custom light curve extraction pipeline starting from Simple Aperture Photometry (SAP; Twicken et al., 2010; Morris et al., 2020) fluxes from the Science Processing Operations Center (SPOC; Jenkins et al., 2016). We used the shortest cadence data available for each sector. All TESS light curves used in this analysis can be found in MAST: http://dx.doi.org/10.17909/9deq-2151 (catalog 10.17909/9deq-2151)
We applied systematic corrections following the prescription in Vanderburg et al. (2019). To summarize, we corrected the SPOC SAP light curves with a linear model consisting of a basis spline (B-spline) with regularly-spaced breaks at 0.2 day intervals to model long-term, low-frequency stellar variability, several moments of the distribution of the spacecraft quaternion time series measurements within each exposure, seven co-trending basis vectors from the SPOC Presearch Data Conditioning (PDC; Stumpe et al., 2012; Smith et al., 2012; Stumpe et al., 2014) band-3 flux time series correction with the largest eigenvalues, and a high-pass (0.1 day) time series from the SPOC background aperture. Additional details required to reproduce this correction can be found in Vanderburg et al. (2019).
@@ 2.1 Extraction pipeline
3 Identification of the transit
As part of our search for transiting planets in the youngest stellar associations (Barber et al. submitted), and motivated by both the single-transit detection in Rizzuto et al. (2020) and the recent transit timing variations (TTVs) detected in HIP 67522 b (Thao et al. submitted), we searched HIP 67522 for additional planet signals. We used the updated N&L pipeline (Rizzuto et al., 2017)111https://github.com/arizzuto/Notch_and_LOCoR as described in Barber et al. (submitted). To summarize, we updated the default box-least squares (BLS) search to the more recent implementation in astropy222http://www.astropy.org (version 4.2; Astropy Collaboration et al., 2018) and used a more optimized period and duration search grid.
@@ 3 Identification of the transit
4 Ground-based follow up
4.1 MuSCAT4
@@ 4.2 Sinistro
5 Stellar Properties
Rizzuto et al. (2020) estimated the stellar radius (), effective temperature (= K), and luminosity () of HIP 67522 by fitting the spectral-energy-distribution with a grid of templates. They estimated the stellar mass () and age by interpolating the observed properties onto the PARSECv1.2S (Bressan et al., 2012) and BHAC15 (Baraffe et al., 2015) stellar evolutionary models.
Since the discovery, there have been new analyses of the Sco-Cen region that may impact these results. Ratzenböck et al. (2023) found the host star was in the -Cen group, which they assign an age between 9.5 and 15.7 Myr depending on the model and input photometry, consistent with the Rizzuto et al. (2020) estimate (172 Myr). Kerr et al. (2021) placed HIP 67522 in the ‘unclustered’ region of Upper Centarus Lupus and were unable to assign a more precise age.
We also repeated the analysis from Rizzuto et al. (2020) using the Gaia DR3 parallax and photometry (Gaia Collaboration et al., 2023). All parameters were within 1 of the values from Rizzuto et al. (2020). Because these and the above age determinations would result in negligible changes, we opted to keep the stellar parameters from Rizzuto et al. (2020) for all analyses.
6 Planet properties
6.1 TESS transit fit
We fit the systematics-corrected TESS using MISTTBORN (MCMC Interface for Synthesis of Transits, Tomography, Binaries, and Others of a Relevant Nature; Mann et al., 2016; Johnson et al., 2018), which uses BATMAN to generate model transits (Kreidberg, 2015), celerite to model stellar variability with a Gaussian Process (GP; Foreman-Mackey et al., 2017), and emcee to explore the parameter space (Foreman-Mackey et al., 2013). Prior to fitting, we manually removed a large flare from 2460059.595 to 2460059.7 BJD.
We fit both planets (b and c) and the stellar variability simultaneously, with 19 fit parameters in total. For each planet, we fit for the time of periastron (), planet orbital period (), planet-to-star radius ratio (), impact parameter (), and and to account for orbital eccentricity () and the argument of periastron (). We fit both planets with a common stellar density () and quadratic limb-darkening coefficients following triangular sampling (, ; Kipping, 2013).
The remaining four parameters were used to fit for the stellar variability in the GP model. We used a stochastically-driven damped simple harmonic oscillator (SHO), following Gilbert et al. (2022), with a jitter term. The free parameters were the undampened oscillator period (), the standard deviation of the process (), the damping timescale (), and the jitter term (). This is a modified version of the RotationTerm in Celerite2 (Foreman-Mackey, 2018).
Most parameters evolved under uniform priors with only physical limitations. The TTV amplitude is small ( minutes), with the TESS data alone being consistent with a linear ephemeris (Thao et al. submitted), and did not impact our results. For , we used a Gaussian prior from the SED and isochrone fits reported in Rizzuto et al. (2020), / = . With this prior on , the transits primarily constrain eccentricity and , instead of (see Dawson & Johnson, 2012; Van Eylen & Albrecht, 2015). We applied a Gaussian prior on the limb-darkening coefficients ( and ) derived using LDTK toolkit (Parviainen & Aigrain, 2015) and our adopted stellar parameters.
@@ 6.1 TESS transit fit
6.2 Ground-based transit fits
@@
In the MuSCAT4 data, there is a perturbation between 2460422.985 and 2460423.05 BJD. Similarly, in the Sinistro simultaneous transits, all three datasets exhibit a small spike in flux around 2460437.445 BJD. As we discuss further in Section 8, these were most likely spot crossings and were also seen in the TESS data. Some events may be flares (particularly the first bump in the MuSCAT4 data), however, our Gaussian spot model (below) described the deviations well.
@@
7 False positive analysis
@@
The spot-crossings indicate the transiting/eclipsing body is passing in front of an active star. This favors HIP 67522, as an unassociated background/foreground star is unlikely to be so active.
@@
8 Summary and Conclusions
We report the discovery and validation of HIP 67522 c, originally identified as a single transit event by Rizzuto et al. (2020). We find HIP 67522 c to be 0.71 (7.9 ) and near the 2:1 mean motion resonance (MMR) with HIP 67522 b. We confirm the period and transit depth with two follow-up ground-based transit observations taken with MuSCAT4 and Sinistro. The HIP 67522 system is now the youngest known transiting multi-planet system to-date.
@@ 8 Summary and Conclusions
Many of the young transiting multi-planet systems are also near MMR (e.g. AU Mic b & c (9:4) (Plavchan et al., 2020; Martioli et al., 2021), V1289 Tau c & d (3:2) and d & b (2:1) (David et al., 2019b, a; Feinstein et al., 2022), HD 109833 b & c (3:2) (Wood et al., 2023), and TIC 434398831 b & c (5:3) (Vach et al., 2024a, b). Simulations suggest that planets will form in MMR and will begin to drift once the disk dissipates (5 Myr) (Izidoro et al., 2017), although many mature planets are found at or near MMR (e.g. Steffen & Hwang, 2015). Hamer & Schlaufman (2024) suggest that MMRs should be most common in systems from 10-100 Myr, and Dai et al. (2024) note that such an excess is visible in the known population of young planets.
@@
HIP 67522 b was a Cycle 1 JWST target (Mann et al., 2021). The resulting transmission spectrum showed strong features, with the transit depth varying by 50% in the CO2 and H2O bands, consistent with a low-density planet (Thao et al. submitted). HIP 67522 c is only 20% smaller than HIP 67522 b (0.71 vs 0.89). Assuming it also has a low density, as predicted for such young planets (Owen, 2020), it will be a similarly high-priority target for JWST.
@@ Acknowledgments
The authors wish to thank Halee, Wally, Penny, and Bandit for their invaluable support. M.G.B. was supported by the NSF Graduate Research Fellowship (DGE-2040435), the NC Space Grant Graduate Research Fellowship, and the TESS Guest Investigator Cycle 5 program (21-TESS21-0016; NASA Grant number 80NSSC24K0880). P.C.T. was supported by the NSF Graduate Research Fellowship (DGE-1650116), the Jack Kent Cooke Foundation Graduate Scholarship, and the JWST GO program (2498). A.W.M. was supported by the NSF CAREER program (AST-2143763) and a grant from NASA’s Exoplanet Reseach Program (XRP 80NSSC21K0393). This work is partly supported by JSPS KAKENHI Grant Number JPJP24H00017 and JSPS Bilateral Program Number JPJSBP120249910.